The ISS in the orbit 408 km away from the surface of the earth. Calculate the time taken to orbit the earth.
1
Expert's answer
2021-07-23T12:51:08-0400

According to the third Kepler's law, the time taken to orbit the earth "T" and the orbit radius "a" are connected as follows (see https://en.wikipedia.org/wiki/Kepler%27s_laws_of_planetary_motion#Third_law):


"\dfrac{a^3}{T^2}\approx \dfrac{GM}{4\pi^2}"

where "G = 6.674\times 10^{-11}m^3\cdot kg^{-1}\cdot s^{-2}" is the gravitational constant, and "M = 5.974\times 10^{24} kg" is the mass of the Earth.

The radius of the orbit is the elevation above the surface plus the radius of the Earth itself: "a = 408km +6371km =6.779\times 10^6m".

Expressing "T", obtain:


"T = \sqrt{\dfrac{4\pi^2a^3}{GM}} = 2\pi\sqrt{\dfrac{a^3}{GM}}\\
T = 2\pi\sqrt{\dfrac{(6.779\times 10^6)^3}{6.674\times 10^{-11}\cdot 5.974\times 10^{24}}} \approx 5554s"

Answer. 5554 s.